# Dual comb–enhanced microwave clock synchronization over commercial fiber

### v1

@@ III Conclusion
In summary, we have demonstrated a femtosecond-level time synchronization between remote microwave clock–based timescales over a commercial fiber link of 205.86 km. The residual time deviation after synchronization was 6.23 fs at 1 s, and the instability was below for an averaging time longer than 10,000 s with the dominate noise from the microwave clock–based LOS measurement. This configuration can make the high time-resolution feature of the optical frequency comb compatible with most existing fiber infrastructures, which greatly promotes the application range of fiber-based time-frequency networks. This result fully merges the bottleneck of time synchronization based on microwave clocks in a hybrid time-frequency network to achieving the interconnection between any microwave and optical clocks. Thus, this method has the potential for microwave-based precise applications such as positioning, navigation, timing and coherent sensing. Moreover, this technique enhances the probe ability of subtle variations of fiber, which will generate new possibilities for channel modeling and a secure time-transfer network to resist potential delay attacks. An interesting extension of this work will further investigate the quantum-limited properties [12] in the microwave clock–based fiber scenarios.
IV Methods
@@
Asynchronous locking system. In our experiment, the repetition frequencies of two combs were phase-locked to a microwave clock to guarantee the coherence of the pulse-envelope phase, while the CEO frequency was free running. The reason is that the CEO frequency has negligible influence on the envelope measurement at an appropriate sampling rate [34].
@@
To maintain the pulse-envelope phase synchronization between two combs with a small frequency offset, we developed an asynchronous locking system, which contained a high-precision single-input two-output DDS, and two fiber-loop optical-microwave phase detectors (FLOM-PDs) [51]. The single-input two-output DDS was used to generate two electrical signals with repetition frequencies of 1 GHz and 1.00001 GHz, which were synchronized to the 10-MHz microwave clock and served as the references of two combs. To maintain the synchronization between the outputs of the DDS and two combs, we locked the repetition frequencies of two combs to the outputs of the DDS using the FLOM-PD, where the 10-th harmonic of the respective combs were used to improve the locking stability. The overall locking-induced time deviation of the system was 45 fs at 1 s.
@@
Eliminating of the ambiguity. Instead of the modulated laser scheme to remove the ambiguity range, we used the MZM to directly encode time markers on the intensity of the optical pulses. The MZM was driven by the digital-to-analog converters (DAC) synchronized to the local clock. High control voltage was supplied to suppress the intensity of the optical pulse to label data “0”, whereas no control voltage was supplied to maintain the intensity of the optical pulses to label data “1”.
@@
Out-of-loop verification. To verify the time offset of the clocks after synchronization, the out-of-loop measurement of the time offset , which is independent from the in-loop calculated time offset , is necessary. Here, we also filtered an output on both combs at each site to generate the out-of-loop reference interferogram that is independent from the in-loop signals to verify the synchronization results. The out-of-loop reference interferograms at two sites contained the timing information of two clocks and could directly eliminate uncertainty from the photoelectronic detection. Then, the out-of-loop verification was performed by measuring the relative time interval of two out-of-loop reference interferograms from two sites using an ADC and transmitted to an computer with a real-time data refresh rate of 1 s, which is independent of the FPGA processor used for synchronization. Therefore, the PSD is available only for Fourier frequencies below 0.5 Hz.
V DATA AVAILABILITY
The data that support the findings of this study are available from the corresponding author upon reasonable request.
VI CODE AVAILABILITY
The code used in this study is available from the corresponding author upon reasonable request.

### v2

@@ V Conclusion
In summary, we have demonstrated a femtosecond-level time synchronization between remote microwave clock–based timescales over a commercial fiber link of 205.86 km. The residual time deviation after synchronization was 6.23 fs at 1 s, and the instability was below for an averaging time longer than 10,000 s with the dominant noise from the microwave clock–based LOS measurement. This configuration can make the high time-resolution feature of the optical frequency comb compatible with most existing fiber infrastructures, which greatly promotes the application range of fiber-based time-frequency networks. This result fully merges the bottleneck of time synchronization based on microwave clocks in a hybrid time-frequency network to achieving the interconnection between any microwave and optical clocks. Thus, this method has the potential for microwave-based precise applications such as positioning, navigation, timing and coherent sensing. Moreover, this technique enhances the probe ability of subtle variations of fiber, which will generate new possibilities for channel modeling and a secure time-transfer network to resist potential delay attacks. An interesting extension of this work will further investigate the quantum-limited properties [12] in the microwave clock–based fiber scenarios.
VI Methods
VI.1 Asynchronous locking system
@@
In our experiment, the repetition frequencies of two combs were phase-locked to a microwave clock to guarantee the coherence of the pulse-envelope phase, while the CEO frequency was free running. The reason is that the CEO frequency has negligible influence on the envelope measurement at an appropriate sampling rate [34].
@@
To maintain the pulse-envelope phase synchronization between two combs with a small frequency offset, we developed an asynchronous locking system, which contained a high-precision single-input two-output DDS, and two fiber-loop optical-microwave phase detectors (FLOM-PDs) [54]. The single-input two-output DDS was used to generate two electrical signals with repetition frequencies of 1 GHz and 1.00001 GHz, which were synchronized to the 10-MHz microwave clock and served as the references of two combs. To maintain the synchronization between the outputs of the DDS and two combs, we locked the repetition frequencies of two combs to the outputs of the DDS using the FLOM-PD, where the 10-th harmonic of the respective combs were used to improve the locking stability. The overall locking-induced time deviation of the system was 45 fs at 1 s, and the corresponding phase noise at typical frequency points were -55 , -100 , -123 , and the floor of -130 was obtained. Future improvements in the FLOM-PD performance can be attained through the use of high-order harmonics for locking and optimizing the noise characteristics of electronic devices.
VI.2 Elimination of the ambiguity
@@
Instead of the modulated laser scheme to remove the ambiguity range, we used the MZM to directly encode time markers on the intensity of the optical pulses. The MZM was driven by the digital-to-analog converters (DAC) synchronized to the local clock. High control voltage was supplied to suppress the intensity of the optical pulse to label data “0”, whereas no control voltage was supplied to maintain the intensity of the optical pulses to label data “1”.
@@
VI.3 Out-of-loop verification
To verify the time offset of the clocks after synchronization, the out-of-loop measurement of the time offset , which is independent from the in-loop calculated time offset , is necessary. Here, we also filtered an output on both combs at each site to generate the out-of-loop reference interferogram that is independent from the in-loop signals to verify the synchronization results. The out-of-loop reference interferograms at two sites contained the timing information of two clocks and could directly eliminate uncertainty from the photoelectronic detection. Then, the out-of-loop verification was performed by measuring the relative time interval of two out-of-loop reference interferograms from two sites using an ADC and transmitted to a computer with a real-time data refresh rate of 1 s, which is independent of the FPGA processor used for synchronization. Therefore, the PSD is available only for Fourier frequencies below 0.5 Hz. The limited sampling rate hampers the acquisition of PSD spectra at high Fourier frequencies, thereby posing difficulties in estimating MDEVs using the integration method described in Refs. [55, 56]. In the future, we hope to increase the refresh rate of the our out-of-loop verification system to obtain more complete PSD spectrum information.
VII Funding
This work was supported by the National Natural Science Foundation of China (Grants Nos. 62201012), and the National Hi-Tech Research and Development (863) Program.
VIII Author contributions
Z.C., D.Y., G.L. and Y.Z. performed the experiment and overall data analysis. Z.C., D.Y. and Y.Z. performed the theoretical analysis. Z.C., D.Y., and G.L. processed the data. Z.C. and H.G. contributed to all parts of the work. Z.C. and D.Y. conceived the experiment, and Z.C., S.Y. B.L. and H.G. supervised the project. All of the authors were involved in the discussion and interpretations of the results. Z.C. wrote the manuscript, and D.Y., G.L., Y.Z., S.Y., B.L. and H.G. provided revisions.
IX Disclosures
The authors declare no conflicts of interest.
X Data Availability
Data supporting the results presented in this paper are available from the authors upon reasonable request.